# thanks: Brinson Prize Fellow

### v1

@@ VII.1 Comparison with previous work
Our results are broadly consistent with previous inferences from gravitational imaging [Vegetti++14, Birrer++17, Hezaveh++16, Ritondale++19], although the uncertainties from imaging analyses are significantly larger than those presented in this work. The larger uncertainties mainly result from extrapolating individual halo detections to statements on the full halo population. This complication aside, results from gravitational imaging offer a complementary tool to identify aspects of halo substructure that could impact population-level constraints. For example, the recent detection of an extremely compact $\sim 10^{6}M_{\odot}$ object in an extended arc [Powell++25] suggests the existence of a population of similar objects. Assuming the perturber is not a globular cluster (which we already include in our model), if the number density of such objects is high enough, we would detect their presence in the flux ratio statistics of lensed quasars.
@@ VII.2 Interpretation
Our constraints on the free-streaming length are stronger than those obtained from other small-scale probes, including previous lensing studies [Birrer++17, Gilman++20, Hsueh++20, Keeley++24], the Lyman-$\alpha$ forest [Villasenor++23, Irsic++24], dwarf galaxies [Newton++21, Nadler++21, Dekker++22, Tan++25] and stellar streams [Banik++21, Nibauer++25]. Our limits are to be considered robust as the numerous improvements to our modeling framework overall weaken the bounds on WDM for a given dataset with respect to previous analyses. Our measurement of the projected mass density in CDM subhalos is consistent with the predictions from the semi-analytic model galacticus to within $2\sigma$ [Gannon++25, Du++25]. It is in mild tension (inferring a higher value) with the predictions from the recent Symphony suite of $N$-body simulations [Nadler++23, Gannon++25].
@@ Acknowledgments
Part of this work was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA.
@@ Appendix C Globular clusters
Figure 22: The joint distribution of flux ratios for a mock lens with a cusp image configuration that maximizes sensitivity to perturbations by halos and globular clusters. Contours show the flux-ratio distributions in CDM (black), WDM (blue), both including globular clusters with $\Sigma_{\rm{gc}}=5\times 10^{5}\mathrm{M}_{\odot}\ \rm{kpc^{-1}}$ and $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$. The magenta and red distributions include perturbations by globular clusters only, and the yellow shows the prediction of a smooth lens model, without halos or GCs. The inset shows a schematic of the mock lens image configuration and magnifications used for the calculation.
@@
We can quantify the lensing signal associated with a population of globular clusters by examining how different assumptions regarding the GC mass functions and projected number density affect flux ratio statistics. We make this comparison in relation to the effect of dark matter halos in CDM and WDM to illustrate the relative importance of injecting GCs in the lens model, and changing the properties of dark matter halos. In Figure 22 we show the joint distribution of flux ratios predicted for a mock lens system with a cusp image configuration, which maximizes the sensitivity to small-scale perturbations. The black distribution corresponds to CDM subhalos and field halos only (no GCs) with $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-1}}$. The blue distribution shows the WDM prediction, assuming $m_{\rm{hm}}=10^{8}\mathrm{M}_{\odot}$ and $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-1}}$. The magenta distribution shows perturbations by globular clusters with a surface mass density $\Sigma_{\rm{gc}}=10^{6}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ and an average GC mass of $\langle m_{\rm{gc}}\rangle=5\times 10^{5}\mathrm{M}_{\odot}$, and the red distribution has $\Sigma_{\rm{gc}}=2\times 10^{5}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ and $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$. For reference, in our analysis we assumed $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$ and $\Sigma_{\rm{gc}}=10^{5.6}\mathrm{M}_{\odot}$. The width of each log-normal mass function is 0.6 dex. The yellow distribution shows the model-predicted flux ratios with no perturbation from GCs or halos.
@@ Appendix D Baseline lens models
Notebooks that reproduce the baseline lens models, shown in Figures 23, 24, 25, 26, and 27, are publicly available  [^29] https://github.com/dangilman/samana/tree/main/notebooks. The source light model complexity used in the baseline model follows from the calculation of $n_{\rm{max}}$ that minimizes the BIC, or for which the model-predicted flux ratios converges, as discussed in Section II.2.3.
Table 3: The effective macromodel parameters inferred for each lens, after marginalizing over substructure in the lens model. We refer to these as effective parameters because they will absorb some features of the (sub)halo population. Therefore, uncertainties will be larger than those typically obtained from strong lens modeling without substructure. We quote results only for the parameters constrained by the imaging data importance weights $w_{\rm{img}}$, full lens model reconstructions can be found in the samana code repository. For systems without imaging data (PSJ0147, RXJ0911, WFI2026, B1422) we quote the parameters constrained using image positions and flux ratios. From left: Lens system, main deflector mass profile normalization, main deflector axis ratio, external shear strength, logarithmic slope of main deflector mass profile, $m=3$ and $m=4$ multipole strengths (the $m=1$ term is unconstrained for most systems), and the mass of the nearest satellite or companion galaxy (only WFI2033 has more than one satellite/companion). Uncertainties are $68\%$ confidence intervals. Cases where one of the multipole strengths has a median $\approx 0$ but large uncertainties (e.g. $a_{4}$ in J0405, J1042) indicate the distribution is bimodal.

### v2

@@ VII.1 Comparison with previous work
Our results are broadly consistent with previous inferences from gravitational imaging [Vegetti++14, Birrer++17, Hezaveh++16, Ritondale++19], although the uncertainties from imaging analyses are significantly larger than those presented in this work. The larger uncertainties mainly result from extrapolating individual halo detections to statements on the full halo population. This complication aside, results from gravitational imaging offer a complementary tool to identify aspects of halo substructure that could impact population-level constraints. For example, the recent detection of an extremely compact $\sim 10^{6}M_{\odot}$ object in an extended arc [Powell++25] could suggest the existence of a population of similar objects. Assuming the perturber is not a globular cluster (which we already include in our model), if the number density of such objects is high enough, we may detect their presence in the flux ratio statistics of lensed quasars.
@@ VII.2 Interpretation
Our constraints on the free-streaming length are stronger than those obtained from other individual small-scale probes, including previous lensing studies [Birrer++17, Gilman++20, Hsueh++20, Keeley++24], the Lyman-$\alpha$ forest [Villasenor++23, Irsic++24], dwarf galaxies [Newton++21, Nadler++21, Dekker++22, Tan++25] and stellar streams [Banik++21, Nibauer++25], although probe combination between lensing and dwarf galaxies has resulted in stronger constraints [Nadler++21]. Our limits are to be considered robust, as the numerous improvements to our modeling framework overall weaken the bounds on WDM for a given dataset with respect to previous analyses. Our measurement of the projected mass density in CDM subhalos is consistent with the predictions from the semi-analytic model galacticus to within $2\sigma$ [Gannon++25, Du++25]. It is in mild tension (inferring a higher value) with the predictions from the recent Symphony suite of $N$-body simulations [Nadler++23, Gannon++25].
@@ Acknowledgments
The work of LAM and DS was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004).
DS acknowledges the support of the Fonds de la Recherche Scientifique-FNRS, Belgium, under grant No. 4.4503.1 and the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) under contract number 4000142531.
@@
This research is based, in part, on data collected at the Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. We are honored and grateful for the opportunity of observing the Universe from Maunakea, which has cultural, historical, and natural significance in Hawaii.
@@ Appendix C Globular clusters
Figure 22: The joint distribution of flux ratios for a mock lens with a cusp image configuration that maximizes sensitivity to perturbations by halos and globular clusters. Contours show the flux-ratio distributions in CDM (black), WDM (blue), both including globular clusters with $\Sigma_{\rm{gc}}=5\times 10^{5}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ and $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$. The magenta and red distributions include perturbations by globular clusters only, and the yellow shows the prediction of a smooth lens model, without halos or GCs. The inset shows a schematic of the mock lens image configuration and magnifications used for the calculation.
@@
We can quantify the lensing signal associated with a population of globular clusters by examining how different assumptions regarding the GC mass functions and projected number density affect flux ratio statistics. We make this comparison in relation to the effect of dark matter halos in CDM and WDM to illustrate the relative importance of injecting GCs in the lens model, and changing the properties of dark matter halos. In Figure 22 we show the joint distribution of flux ratios predicted for a mock lens system with a cusp image configuration, which maximizes the sensitivity to small-scale perturbations. The black distribution corresponds to CDM subhalos and field halos only (no GCs) with $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-1}}$. The blue distribution shows the WDM prediction, assuming $m_{\rm{hm}}=10^{8}\mathrm{M}_{\odot}$ and $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-1}}$. The magenta distribution shows perturbations by globular clusters with a surface mass density $\Sigma_{\rm{gc}}=10^{6}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ and an average GC mass of $\langle m_{\rm{gc}}\rangle=5\times 10^{5}\mathrm{M}_{\odot}$, and the red distribution has $\Sigma_{\rm{gc}}=2\times 10^{5}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ and $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$. For reference, in our analysis we assumed $\langle m_{\rm{gc}}\rangle=2\times 10^{5}\mathrm{M}_{\odot}$ and $\Sigma_{\rm{gc}}=10^{5.6}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$. The width of each log-normal mass function is 0.6 dex. The yellow distribution shows the model-predicted flux ratios with no perturbation from GCs or halos.
@@ Appendix D Baseline lens models
Notebooks that reproduce the baseline lens models, shown in Figures 23, 24, 25, 26, and 27, are publicly available  [^29] https://github.com/dangilman/samana/tree/main/notebooks. The source light model complexity used in the baseline model follows from the calculation of $n_{\rm{max}}$ that minimizes the BIC, or for which the model-predicted flux ratios converges, as discussed in Section II.2.3. Table 3 lists the macromodel parameters inferred for each lens using the image positions, flux ratios, and lensed arcs.
Table 3: The effective macromodel parameters inferred for each lens, after marginalizing over substructure in the lens model. We refer to these as effective parameters because they will absorb some features of the (sub)halo population. Therefore, uncertainties will be larger than those typically obtained from strong lens modeling without substructure. We quote results only for the parameters constrained by the imaging data importance weights $w_{\rm{img}}$, full lens model reconstructions can be found in the samana code repository. For systems without imaging data (PSJ0147, RXJ0911, WFI2026, B1422) we quote the parameters constrained using image positions and flux ratios, for the rest we use image positions, flux ratios, and lensed arcs. From left: Lens system, main deflector mass profile normalization, main deflector axis ratio, external shear strength, logarithmic slope of main deflector mass profile, $m=3$ and $m=4$ multipole strengths (the $m=1$ term is unconstrained for most systems), and the mass of the nearest satellite or companion galaxy (only WFI2033 has more than one satellite/companion). Uncertainties are $68\%$ confidence intervals. Cases where one of the multipole strengths has a median $\approx 0$ but large uncertainties (e.g. $a_{4}$ in J0405, J1042) indicate the distribution is bimodal.